OpenStudy (anonymous):

EASY MEDAL Simplify: square root of 10 times square root of 8 square root of 18 4 square root of 5 square root of 80 None of the above

OpenStudy (gorv):

\[\sqrt{10*\sqrt{8}}\]

OpenStudy (gorv):

is it ur Q???? @ZekeWars

OpenStudy (anonymous):

\[\sqrt{10*2\sqrt{2}}\]

OpenStudy (anonymous):

\[\sqrt{10} * \]

OpenStudy (anonymous):

\[\sqrt{10} * \sqrt{8}\]

OpenStudy (anonymous):

that one

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

@bookworm00981

OpenStudy (anonymous):

@Machida

OpenStudy (anonymous):

@WassupJade

OpenStudy (anonymous):

\[\sqrt{10}* \sqrt{8}\]

OpenStudy (anonymous):

how do i solve that?

OpenStudy (gorv):

\[\sqrt{10*8}=\sqrt{80}=\sqrt{16*5}=4*\sqrt{5}\]

OpenStudy (anonymous):

Sorry, mathematics isn't my strong subject.

OpenStudy (anonymous):

Could you be more clear with your question please?